OpenStudy (anonymous):

Hello, Is a math function always undefined if you divide by zero, even if there is a +3 after? Example: (X/X-3)+3.

terenzreignz (terenzreignz):

Trying to check if this function is defined at X = 3 ?

OpenStudy (anonymous):

Yes, but i dont know if that 3 makes the function defined always.

terenzreignz (terenzreignz):

It's not... lol, hang on, let me explain...

terenzreignz (terenzreignz):

Suppose it IS defined...at x = 3. Then we have a value for it, say k. \[\Large k = \frac30 + 3\] right?

terenzreignz (terenzreignz):

Then we can subtract 3 from both sides \[\Large k -3 = \frac30\] And divide 3 from both sides \[\Large \frac{k-3}3= \frac10\] And suddenly, 1/0 has a value, which we know to be false... thus, \[\Large \frac{x}{x-3}+3\] is undefined, even if there's a + 3

OpenStudy (anonymous):

Thanks, got it!

OpenStudy (skullpatrol):

@terenzreignz where did you get "a + 3" ?

terenzreignz (terenzreignz):

sorry, 'a' here is the English indefinite article, not a variable... I apologize for the ambiguity ^_^

terenzreignz (terenzreignz):

Probably should have typed a +3

terenzreignz (terenzreignz):

or a "+3"

OpenStudy (skullpatrol):

Ahhh..a plus 3.

OpenStudy (skullpatrol):

Nice work.

OpenStudy (ybarrap):

Note also that x = 3/0 is undefined because then we would be able to say that x * 0 = 3, which says that there is something that multiplies with 0 to give a three. But "0" added to itself "x" times is still zero and could never be 3. Zero has this unique property that x*0 = 0 for any x. So x/0 is undefined because x*0=3 is undefined. We even agree to say x=0/0 is undefined, although x*0 = 0 is defined. So anything divided by zero, stops us in our tracks and we are prohibited from using that value.